Theresa has $1.75 in quarters and nickles. If she has 15 coins total, how many quarters does she have?
Answer by jorel1380(3719)

q+n=15
25q+5n=175
5q+5n=75
20q=100
q=5
n=10
Theresa has 10 nickels and 5 dimes..